OpenStudy (anonymous):

please helppppp a bag contains 3 blue, 4 red, and 2 whits marbles. three times you draw a marble and return it. find the probability of getting a red, then white then blue

OpenStudy (anonymous):

First you have to calculate the total number of marbles. 3+4+2. Second you need to divide the number of red marbles by the total number of marbles. 4/(3+4+2). That is the total number of red marbles if you drew 1 marble 1 time. Instead, you need to multiply 4/(3+4+2) by 3. Then you will have the total probability of red marbles you can draw assuming you draw three times. Quick recap. # of marbles of one color divided by the total number of marbles. All of that times three. Let me know if you have any questions.

OpenStudy (anonymous):

\[\frac{4}{9}\times \frac{2}{9}\times \frac{3}{9}\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yw